OpenStudy (anonymous):

If someone put $4000 in a savings account.At the end of 3 years,the account had earned$960 in simple interest. At what annual simple intrest rate did the account grow explain? How many more dollars would she have in the account if the intrest rate was 1% greater?

OpenStudy (amistre64):

use the formula: I = Prt, and solve for r

OpenStudy (anonymous):

I know the formula but I don't know how to use the formula to find r...

OpenStudy (amistre64):

plug in whats given ... I is interest earned, P is starting principal, and t is time spent

OpenStudy (amistre64):

960 = 4000(3)r ... how do we find r now?

OpenStudy (anonymous):

i dunno

OpenStudy (anonymous):

I'm really bad at math

OpenStudy (amistre64):

you have to have some basic idea what to do ... im sure this question doesnt exist ina vaccuum. any idea you have would be helpful to work from

OpenStudy (anonymous):

Divide maybe?

OpenStudy (anonymous):

I know how to do simple interest but not backwards

OpenStudy (amistre64):

division is correct :) so, we have 960 = 12000r, dividing gets us: r = 960/12000, and simplify as desired

OpenStudy (amistre64):

then the question asks if the rate is r + .01 since we need to convert 1% to a decimal form

OpenStudy (anonymous):

r=.08 so .08+.01=.09?

OpenStudy (amistre64):

r = .08 is the first part, now use r=.09 to determine the interest for the second part :)

OpenStudy (amistre64):

or rather: r = 8% as is, what if r was 9% how much interest would have been earned?

OpenStudy (anonymous):

okay so...$4000*.09*3?

OpenStudy (amistre64):

yep

OpenStudy (anonymous):

Which would be...1080?

OpenStudy (amistre64):

correct. so the interest earned at 9% would have been 1080. so the solution is either: 1080, or maybe the total in the acowingspan may be 4000+1080. I cant really tell.

OpenStudy (amistre64):

lol, account ... typoed it into a bad word apparently

OpenStudy (anonymous):

Thankyou so much!! Sometimes I just need to be walked through the problem. :)

OpenStudy (anonymous):

It's k

OpenStudy (amistre64):

critical thinking skills are acquired :) good luck

OpenStudy (texaschic101):

I = PRT P = 4000 R = R T = 3 I = 960 now we sub 960 = 4000 * R * 3 960 = 12000R 960/12000 = R .08 = R 8% = R Now it is 8% for 3 years, so it grew at .08/3 = .0266 = 2.7% per year If the interest was 1% greater... I = PRT P = 4000 R = 9% or .09 T = 3 now we sub I = 4000 * .09 * 3 I = 1080 4000 + 1080 = 5080 would be in the account after 3 years at 9% interest